# Learning Machine Morality through Experience and Interaction

### v2

@@ 4 Evaluating Moral Learning Agents
4.1 Outcomes, Behaviors and Qualitative Assessment
After training such morally motivated agents, researchers must evaluate whether their learning was ethically effective. This requires the definition of a set of outcome metrics. Traditional Reinforcement Learning research would train agents in a specific simulated environment and measure some population-level outcome such as cumulative reward. However, as discussed in Section 2.2, further evaluation metrics must be designed to identify potential reward-hacking behavior of RL agents - in other words, we must also measure concepts or behaviors that were not directly encoded in the agents’ reward functions. Ideally, these measurements should be broad enough to cover a vast number of diverse ethical considerations in a wide range of scenarios. The challenges of this have been discussed by ? (?). As an illustration, in Section 4.2 below, we present an example set of population-level outcome metrics that were proposed by ? (?) for the case of social dilemma games.
On top of population-level outcomes, a game-theoretic evaluation of emerging behaviors may also be of interest, evaluating actions taken in the game over the course of learning, and the extent to which these actions are aligned with moral norms. In social dilemmas, this must include a measurement of cooperation displayed by each agent towards the end of the learning period. This allows for a non-consequentialist evaluation of learned behaviors and policies, and provides better insight for norm-based morality.
Finally, an alternative means of evaluation could include a qualitative assessment by a human who interacts with an agent. Within the domain of autonomous vehicles, large-scale evaluations of hypothetical agent behaviors have been conducted by ? (?) using survey studies, and focusing on the trolley problem (?) in particular as an agent environment. Much more qualitative evaluation is required in other domains and sets of problems to understand human perception of AI-assisted decision making more generally.
@@ 4.2 Example: Measuring Outcomes in Social Dilemma Environments
In the dilemma scenarios presented here, the choice of one agent affects the outcomes of another. Thus, it is possible that scenarios will arise in which agents intentionally designed with a certain morality (e.g., maximize ‘collective reward’) actually display behavior that maximizes one outcome at a cost of another (e.g., ‘equality’). For this purpose, on top of individual rewards, researchers have highlighted the importance of measuring a range of social (i.e., population-level) outcomes (?).
The most popular social outcome metric is the collective payoff for all agents - used by ? (?, ?, ?). Additional moral evaluation metrics can include a measure of equality such as the Gini coefficient (?), and an egalitarian measure of the minimum payoff obtained by any agent in the population on every iteration (?). These three social outcome metrics were formalized in ? (?) as a cumulative return $G$ (i.e., the total reward accumulated over all the $T$ iterations of an iterated game) for both players $M$ and $O$ as follows:
@@
In ? (?), the five agents from Table 4 are systematically evaluated on three distinct social dilemma games in terms of the cumulative outcomes defined above and cooperation levels. The results show that - with sufficient exploration and learning time - all types of moral agents are able to learn fully cooperative policies, though the ‘equality’ definition produces the least efficient learning agent, which is reflected in the social outcome metrics.
4.3 Example: Measuring Moral Behavior in Language Models
The domain of Large Language Models (LLMs) provides a different set of challenges for measurement. Due to the open-ended nature of its inputs and outputs, and the likely moral diversity of preferences of its users (discussed above), it is harder to define a standard set of metrics to evaluate the moral learning of LLMs. Nevertheless, a large set of benchmark metrics and data sets have already been proposed for measuring morality in these systems (see ?, ?, ?, ?).
@@ 5 Outlook & Implications
Our discussion suggests that the traditional fully top-down (e.g., rule-based) methods, or the currently popular fully bottom-up approaches to AI morality and safety (e.g., fine tuning of LLMs based on Reinforcement Learning from Human Feedback), whilst intuitively attractive because of the control they seem to impose (top-down), or the generality they may offer (bottom-up), are limited and pose potential risks for systems deployed in the real world. We motivate the use of a combination of interpretable top-down quantitative definitions of moral objectives, based on existing frameworks in fields such as Moral Philosophy, with the bottom-up advantages of trial-and-error learning from experience via RL. We argue that this hybrid methodology provides a powerful way of studying and imposing control on an AI system while enabling adaptation to dynamic environments. We review three case studies combining moral principles with learning, which provide proof-of-concept for the potential of this hybrid approach in creating more prosocial and cooperative agents.
Moreover, we appreciate the fact that quantifying morality by translating complex moral frameworks into elementary reward functions in this way is certainly a simplification of reality. Nevertheless, the original moral philosophical frameworks discussed in this work might be seen as a form of abstraction and, in a sense, simplification of the complex and multi-faceted human condition, and yet these have served the field for centuries.
We believe there is a need to study these hybrid approaches in a variety of social environments for morally-informed learning (such as, for instance, other social dilemma games), possibly involving a large number of players. The specific moral reward functions proposed here and in a number of existing practical works (?, ?, ?, ?), as well the other methodologies reviewed (?, ?), are a starting point for building more sophisticated moral agents, including those based on more than one moral objective (?). In general, the reward-based methods have a disadvantage of the fact that they rely on a clear scalar payoff structure, and in the case of some of the intrinsic moral rewards proposed - a discrete set of actions. An open research question remains about how to obtain this reward structure in open-ended tasks while still maintaining principled control over the morality implemented in the reward.
@@
Finally, it is worth noting that approaches described above can bring benefit back to the disciplines they are founded on and took inspiration from. In particular, the proposed methodology can be used to reflect upon and understand different moral frameworks and their impacts, especially in complex situations. For example, it can be used for simulating the outcomes deriving from the presence in a society of agents with specific moral preferences, or to study emergent behaviors in societies composed of agents with different moral principles. In general, this can be seen as a practical example of ‘Computational Philosophy’ (?). Moreover, in general terms, modeling moral behavior as a reward-maximizing $\epsilon$-greedy choice made by a continually learning agent can provide insights for Evolutionary Game Theory. Our simulations with learning agents can discover how certain moral behaviors can evolve (and might have evolved) in a society given a payoff structure and different distributions of preferences and norms within a population. Additionally, an analysis of the learning dynamics of agents can provide insights for Developmental Psychology, possibly offering a way to test various theories of moral development in children (?, ?).
@@ 6 Conclusion
In this paper, we have made the case for the use of learning with explicitly defined moral principles for developing morality in AI agents, contrasting it with more traditional top-down implementations in AI safety and ethics, and the more recent entirely bottom-up methods for learning implicit values from human feedback, such as those used for fine-tuning Large Language Models. We have suggested RL as an effective mechanism for learning morality from experience in a given environment, and reviewed three case studies which successfully implemented this in agents or the overall training of systems in practice. In one of these case studies, we have outlined a hybrid methodology in which agents learn morality from experience by using intrinsic rewards based on classic moral frameworks, and presented a formalization of a set of specific reward functions using social dilemma environments. Finally, we have proposed an example of potential evaluation of the outcomes and behaviors emerging from interactions between different agents over time.
We hope that this article can motivate further work in the area of learning morality from experience and interaction with the help of top-down explicit moral principles. While there are substantial challenges in implementing the proposed design guidelines in practice, we believe it provides a promising avenue for addressing the growing safety concerns around emerging advanced AI.
@@
Bottom-up Development of a system / inference of values through experience, organically, without planning the general architecture / constraints / components in advance. An example is Reinforcement Learning of desired behaviors from human demonstrations (e.g., via Inverse RL or RLHF).
@@
Moral norm A norm in society about what constitutes moral behavior. An example is ‘not defecting against a cooperator’.
@@ Appendix A. Other Games Involving Morality
Sequential games model social situations in which one player takes an action first, then the other chooses how to respond given how the first player acted. An example of this is the Ultimatum game (?), in which the first player offers a way to split a sum of money, and the second player either accepts the split or rejects it, in which case neither player gets anything. Here the first player may be incentivized to keep a larger sum for themselves, but they risk pushing the opponent too far and ending up with nothing. Applied to ethics, the Ultimatum game can model the principle of fairness in both players (fairness in one’s own behavior for the proposer, and punishment for unfair behavior by the receiver), providing an interesting test bed for moral learning agents.

### v3

@@ 4 Evaluating Moral Learning Agents
4.1 Outcomes, Behaviours and Qualitative Assessment
After training such morally motivated agents, researchers must evaluate whether their learning was ethically effective. This requires the definition of a set of outcome metrics. Traditional Reinforcement Learning research would train agents in a specific simulated environment and measure some population-level outcome such as cumulative reward. However, as discussed in Section 2.2, further evaluation metrics must be designed to identify potential reward-hacking behaviour of RL agents - in other words, one must also measure concepts or behaviours that were not directly encoded in the agents’ reward functions. Ideally, these measurements should be broad enough to cover a vast number of diverse ethical considerations in a wide range of scenarios. The challenges of this have been discussed by ? (?). As an illustration, in Section 4.2 below, we present an example set of population-level outcome metrics that were proposed by ? (?) for the case of social dilemma games. To complement this example of measuring RL agents in social dilemma environments, we also present a discussion of how LLM-based agents can be measured in terms of behavioural alignment (see Section 4.3).
On top of population-level outcomes, a game-theoretic evaluation of emerging behaviours may also be of interest, evaluating actions taken in the game over the course of learning, and the extent to which these actions are aligned with moral norms. In social dilemmas, this must include a measurement of cooperation displayed by each agent towards the end of the learning period as well as the emerging strategies (i.e., state-action combinations chosen by the agent). This allows for a non-consequentialist evaluation of learned behaviours and policies, and provides better insight for norm-based morality.
Finally, an alternative means of evaluation could include a qualitative assessment by a human who interacts with an agent. Within the domain of autonomous vehicles, large-scale evaluations of hypothetical agent behaviours have been conducted by ? (?) using survey studies, and focusing on the trolley problem (?) in particular as an agent environment. Much more qualitative evaluation is required in other domains and sets of problems to understand human perception of AI-assisted decision making more generally.
@@ 4.2 Example: Measuring Outcomes in Social Dilemma Environments
In the dilemma scenarios presented in Section 3.4, the choice of one agent affects the outcomes of another. Thus, it is possible that scenarios will arise in which agents intentionally designed with a certain morality (e.g., maximise ‘collective reward’) actually display behaviour that maximises their desired outcome at a cost of some other important outcome (e.g., ‘equality’). For this purpose, on top of analysing individual rewards, researchers have highlighted the importance of measuring a range of population-level outcomes (?, ?, ?).
The most popular social outcome metric is the collective payoff for all agents - used by ? (?, ?, ?). Additional moral evaluation metrics can include a measure of equality such as the Gini coefficient (?), and an egalitarian measure of the minimum payoff obtained by any agent in the population on every iteration (?). These three social outcome metrics were formalised in ? (?) as a cumulative return $G$ (i.e., the total reward accumulated over all the $T$ iterations of an iterated game) for both players $M$ and $O$ as follows:
@@
In ? (?), the five agents from Table 4 are systematically evaluated on three distinct social dilemma games in terms of the cumulative outcomes defined above and cooperation levels and strategies. The results show that - with sufficient exploration and learning time - all types of moral agents are able to learn fully cooperative policies, though the ‘equality’ definition produces the least efficient learning agent, which is reflected in the social outcome metrics.
4.3 Example: Measuring Moral Behaviour in Language Models
The domain of Large Language Models (LLMs) provides a different set of challenges for measurement. Its inputs and outputs are open-ended, such that the possible action space for a single generated token is as large as the vocabulary size associated with the model. Furthermore, the users of the model are likely to have diverse moral preferences of its users (discussed above). This makes it harder to define a standard set of metrics to evaluate the moral learning of LLMs. Nevertheless, a large set of benchmark metrics and data sets have already been proposed for measuring morality in these systems (see ?, ?, ?, ?). Capitalising on the open-ended nature of human language, studies such as ? (?) have also assessed the generalisation of moral policies from one environment to others, though more remains to be done in terms of assessing their performance in real-world scenarios.
@@ 5 Outlook & Implications
Our discussion suggests that the traditional fully top-down (e.g., rule-based) methods, or the currently popular fully bottom-up approaches to AI morality and safety (e.g., fine tuning of LLMs based on Reinforcement Learning from Human Feedback), whilst intuitively attractive because of the control they seem to impose (top-down), or the generality they may offer (bottom-up), are limited and pose potential risks for systems deployed in the real world. We motivate the use of a combination of interpretable top-down quantitative definitions of moral objectives, based on existing frameworks in fields such as Moral Philosophy, with the bottom-up advantages of trial-and-error learning from experience via RL. We argue that this hybrid methodology provides a powerful way of studying and imposing control on AI systems while enabling adaptation to dynamic environments. We review four case studies combining moral principles with learning, which provide proof-of-concept for the potential of this hybrid approach in creating more prosocial and cooperative agents. We hope that the unifying framework defined in this paper helps future practitioners and policymakers identify new possible avenues for future work.
We appreciate the fact that quantifying morality by translating complex moral frameworks into elementary reward functions, learning constraints or textual principles in this way is certainly a simplification of reality. Nevertheless, the original moral philosophical frameworks discussed in this work might be seen as a form of abstraction and, in a sense, simplification of the complex and multi-faceted human condition, and yet these have served the field for centuries.
We believe there is a need to study these hybrid approaches in a variety of social environments for morally-informed learning, including other social dilemma games and games involving a large number of players. The specific moral reward functions proposed in Tables 3 and 4 and in a number of existing practical works (?, ?, ?, ?), as well the other methodologies reviewed (?, ?), are a starting point for building more sophisticated moral agents, including those based on more than one moral objective (?). In general, the reward-based methods have a disadvantage of the fact that they rely on a clear scalar payoff structure, and in the case of some of the intrinsic moral rewards proposed - a discrete set of actions. An open research question remains about how to obtain this reward structure in open-ended tasks while still maintaining principled control over the morality implemented in the reward.
@@
Finally, it is worth noting that approaches described above can bring benefit back to the disciplines they are founded on and took inspiration from. In particular, the proposed methodology can be used to reflect upon and understand different moral frameworks and their impacts, especially in complex situations. For example, it can be used for simulating human societies, studying the outcomes deriving from the presence of agents with specific moral preferences, or to study emergent behaviours in societies composed of agents with different moral principles. In general, this can be seen as a practical example of ‘Computational Philosophy’ (?). Moreover, in general terms, modelling moral behaviour as a reward-maximising $\epsilon$-greedy choice made by a continually learning agent can provide insights for Evolutionary Game Theory. Our simulations with learning agents can discover how certain moral behaviours can evolve (and might have evolved) in a society given a payoff structure and different distributions of preferences and norms within a population. Additionally, an analysis of the learning dynamics of agents can provide insights for Developmental Psychology, possibly offering a way to test various theories of moral development in children (?, ?, ?).
@@ 6 Conclusion
In this paper, we have made the case for the use of learning with explicitly defined moral principles for developing morality in AI agents, contrasting it with more traditional top-down implementations in AI safety and ethics, and the more recent entirely bottom-up methods for learning implicit values from human feedback, such as those used for fine-tuning LLMs. We have suggested RL as one effective mechanism for learning morality from experience in a given environment, and reviewed a set of case studies to analyse different moral alignment approaches that have been implemented for agentic systems in practice. In two of these case studies (based on the pure-RL or LLM domains), we have outlined a hybrid methodology in which agents learn morality via intrinsic rewards based on classic moral frameworks, and presented a formalisation of a set of specific reward functions using social dilemma environments. We have also reviewed solutions based on alternative approaches, such as constrained RL or fine-tuning LLMs from constitution-based feedback. Finally, we have proposed an example of potential evaluation of the moral outcomes and behaviours emerging from interactions between different agents over time, and outlined the existing open challenged in this area.
We hope that this article can motivate further work in the area of developing moral AI agents via learning with the help of top-down explicit moral principles (i.e., in a hybrid way). While there are substantial challenges in implementing the proposed design guidelines in practice, we believe it provides a promising avenue for addressing the growing safety concerns around emerging advanced AI.
@@
Bottom-up Development of a system / inference of values through experience, organically, without planning the general architecture / constraints / components in advance. An example is Reinforcement Learning of desired behaviours from human demonstrations (e.g., via Inverse RL or RLHF).
@@
Moral norm A norm in society about what constitutes moral behaviour. An example is ‘not defecting against a cooperator’.
@@ Appendix A. Other Games Involving Morality
Sequential games model social situations in which one player takes an action first, then the other chooses how to respond given how the first player acted. An example of this is the Ultimatum game (?), in which the first player offers a way to split a sum of money, and the second player either accepts the split or rejects it, in which case neither player gets anything. Here the first player may be incentivised to keep a larger sum for themselves, but they risk pushing the opponent too far and ending up with nothing. Applied to ethics, the Ultimatum game can model the principle of fairness in both players (fairness in one’s own behaviour for the proposer, and punishment for unfair behaviour by the receiver), providing an interesting test bed for moral learning agents.